[This page intentionally left very nearly blank.] [This page intentionally left very nearly blank.]

Part O Prologue

19

Part I Examples

31 [This page intentionally left very nearly blank.]

Part II Details

43

Part III Conclusion

67 [This page intentionally left very nearly blank.]

Part IV Appendices

75 [This page intentionally left very nearly blank.]

Bibliography

[Abelson et al. 89]

Harold Abelson, Michael Eisenberg, Matthew Halfant, Jacob Katzenelson, Elisha Sacks,

Gerald J. Sussman, Jack Wisdom, and Kenneth Yip

Intelligence in Scientific Computing In CACM, Vol. 32, No. 5, May 1989, pp.546–562

[Abelson, Halfant, Katzenelson & Sussman 88] Harold Abelson, Matthew Halfant, Jacob Katzenelson and Gerald Jay Sussman The Lisp Experience In Annual Reviews in Computer Science, Vol. 3, 1988, pp.167–195 Annual Reviews, Inc., Palo Alto, California, 1998

[Abelson & Sussmans 96] Harold Abelson and Gerald Jay Sussman with Julie Sussman Structure and Interpretation of Computer Programs [2nd Edition] (The MIT Electrical Engineering and Computer Science Series) The MIT Press, 1996

[Adams 91] Stephen Robert Adams Modular Grammars for Programming Language Prototyping Ph.D. Dissertation, University of Southampton, March 1991

Internet URL: http://www-swiss.ai.mit.edu/~adams/MAG/index.html

[Adams & Hanson 93] Stephen Adams and Chris Hanson MIT Scheme User’s Manual Unpublished manuscript, available with the MIT Scheme distribution.

Internet URL: http://www.gnu.org/software/mit-scheme/

[Agarwal, Chaiken & Kranz 91] Anant Agarwal, David Chaiken and David Kranz Multiprocessor Address Tracing and Performance Analysis MIT/LCS/TM-446, MIT Laboratory for Computer Science, Apr 1991 [Agesen & Hölzle 95]

Ole Agesen and Urs Hölzle Type Feedback vs. Concrete Type Inference: A Comparison of Optimisation Techniques

for Object-Oriented Languages

In Proc. ACM OOPSLA ’95, 1995, pp.91–107

[Aho, Hopcroft & Ullman 74] Alfred V. Aho, John E. Hopcroft, and Jeffrey D. Ullman The Design and Analysis of Computer Algorithms Addison-Wesley, 1974

[Aho & Ullman 72] Alfred V. Aho and Jeffrey D. Ullman The Theory of Parsing, Translation, and Compiling, (Volume I: Parsing) (Prentice-Hall Series in Automatic Computation) Prentice-Hall, Inc., 1972

[Aho & Ullman 73] Alfred V. Aho and Jeffrey D. Ullman The Theory of Parsing, Translation, and Compiling, (Volume II: Compiling) (Prentice-Hall Series in Automatic Computation) Prentice-Hall, Inc., 1973

[Aho & Ullman 86] Alfred V. Aho and Jeffrey D. Ullman Compilers: Principles, Techniques and Tools Addison-Wesley, 1986

[Appel, Duba & MacQueen 88] Andrew W. Appel, Bruce F. Duba, and David B. MacQueen Profiling in the Presence of Optimization and Garbage Collection CS-TR-197-88, Princeton University Department of Computer Science, 1988

[Apple 92] Apple Computer, Eastern Research and Technology Dylan: An Object-Oriented Dynamic Language Apple Computer, Inc, 1992

[ARPA 95] Advanced Research Projects Agency, Information Technology Office (ARPA ITO) Evolutionary Design of Complex Software (EDCS)— BAA-9540 (ARPA ITO sponsored research solicitation, broad-area announcement)

Archival URL: http://www.swiss.csail.mit.edu/~mrb/CBD_9540.html
Original URL: http://www.ito.arpa.mil/Solicitations/CBD_9540.html

Close Date: 4:00 PM, EST, Wednesday, November 17, 1995 [Ashley & Consel 94]

J. Michael Ashley and Charles Consel Fixpoint Computation for Polyvariant Static Analyses of Higher-Order Applicative Programs In ACM TOPLAS, Vol. 16, No. 5, Sep 1994, pp.1431–1448

Internet URL: http://repository.readscheme.org/ftp/papers/jmashley/toplas94.pdf

[Asimov 51/52/53,74/82,83/86/90/92] Isaac Asimov The Foundation Trilogy: Foundation, Foundation and Empire, and Second Foundation Gnome Press, 1951, 1952 and 1953 (respectively)

Reprinted as:

The Foundation Trilogy: Three Classics of Science Fiction Avon, 1974 Doubleday & Co., 1982

Reprinted as:

The Foundation Trilogy

Ballantine Books, Inc., 1983, 1986

Doubleday Direct, 1990

Bantam Doubleday Dell Pub., 1992

[Backus 59] J. W. Backus The Syntax and Semantics of the Proposed International Algebraic Language

of the Zurich ACM-GAMM Conference

In Proc. UNESCO International Conf. on Info. Processing, Paris, Jun 1959, pp.125–132

[Backus 78] John Backus Can Programming be Liberated from the von Neumann Style? A Functional Style and Its Algebra of Programs In CACM, Vol. 21, No. 8, Aug 1978, pp.613–641

[Baier & Sager 76] Jean-Loup Baier and Gary R. Sager Dynamic Improvement of Locality in Virtual Memory Systems In IEEE Transactions on Software Engineering, Vol. SE2, No. 1, Mar 1976, pp.54–62

[de Bakker 80] Jaco de Bakker Mathematical Theory of Program Correctness (Prentice-Hall International Series in Computer Science) C. A. R. Hoare (Series Editor) Prentice-Hall International, Inc., 1980 [Ball & Larus 92]

Thomas Ball and James R. Larus Optimally Profiling and Tracing Programs In Proc. ACM POPL ’92, Aug 1992, pp.59–70

[Barendregt 84] Hendrik P. Barendregt The Lambda Calculus: Its Syntax and Semantics [Revised Edition] (Studies in Logic and Foundations of Mathematics, Vol. 103) Elsevier Science Publishing Company, 1984

[Becket 53/55] Samuel Beckett ⟨⟨En attendant Godot^{⟩⟩} Première: Théâtre de Babylone, Paris, France, January 5, 1953 Paris: Editions de Minuit, 1952 (First published trade edition)

Translated as:

Waiting for Godot, a tragicomedy in two acts Première: Arts Theatre, London, England, August 3, 1955 New York: Grove Press, 1954

Internet URL(s): http://samuel-beckett.net/Waiting_for_Godot_Part1.html
                 http://samuel-beckett.net/Waiting_for_Godot_Part2.html

[Beckman & Kamin 98] Mattox Beckman and Sam Kamin Controlled Self-Applicable On-Line Partial Evaluation, Using Strategies In Proc. IEEE Int. Conf. on Computer Languages (ICCL ’98), May 1998, pp.143–152

Internet URL: http://www-sal.cs.uiuc.edu/~kamin/pubs/beckman-iccl98.ps

[Bell 73] James R. Bell Threaded Code In CACM, Vol. 16, No. 6, Jun 1973, pp.370–372

[Bender 2006] Jim Bender Online Bibliography of Partial Evaluation Research

Internet URL:  http://www.partial-eval.org/index.html
E-mail address: mailto:jim@benderweb.net
Futurama Fnord: “Please insert girder.”

[Bentley 82] Jon Louis Bentley Writing Efficient Programs Prentice-Hall, Inc., 1982

[Bentley 86] Jon Bentley Programming Pearls Addison-Wesley Publishing, 1986

[Bentley 89] Jon Bentley More Programming Pearls Addison-Wesley Publishing, 1990

[Berger & Schwichtenberg 91] Ulrich Berger and Helmut Schwichtenberg An Inverse of the Evaluation Functional for Typed λ-calculus In Proc. IEEE Symp. on Logic in Computer Science (LICS ’91), Jul 1991, pp.203–211

Internet URL: (Draft Copy)

   http://www.mathematik.uni-muenchen.de/~berger/articles/lics/inveval.dvi.Z

[Berlin 89] Andrew A. Berlin A Compilation Strategy for Numerical Programs Based on Partial Evaluation S.M. Thesis, MIT AI-TR-1144, MIT Artificial Intelligence Laboratory, Jul 1989

[Berlin 90] Andrew A. Berlin Partial Evaluation Applied to Numerical Computation In Proc. ACM Conference on Lisp and Functional Programming, Jun 1990, pp.139–150

[Berlin & Weise 89] Andrew Berlin and Daniel Weise Compiling Scientific Code using Partial Evaluation MIT AI-Memo-1145, MIT Artificial Intelligence Laboratory, Jul 1989

[Bernoulli 1713,2006] Jakob D. Bernoulli Ars Conjectandi Thurneisser Brothers Press, Basel, 1713

Translated as:

Edith Dudley Sylla (Translator, with Introduction and Notes)

The Art of Conjecturing, together with “Letter to a Friend on Sets in Court Tennis” Johns Hopkins University Press, 2006 [Bjørner, Ershov & Jones 88]

Dines Bjørner, Andrei P. Ershov and Neil D. Jones (Editors) Partial Evaluation and Mixed Computation Proc. IFIP TC2 Workshop on Partial Evaluation and Mixed Computation

Gammel Avernæs, Ebberup, Denmark, Oct 18–24, 1987

Amsterdam: North-Holland, 1988

Internet URL: http://partial-eval.org/events/pemc1987.html

[Blair 90] Michael R. Blair FIX: A Design for an Efficient Type/Region/Effect/Kind Checking System S.M. Thesis, MIT Laboratory for Computer Science, Sep 1990

Internet URL: http://www.swiss.csail.mit.edu/~mrb/papers.html

[Blair 91] Michael R. Blair Fast Structural Type Equivalence and Inclusion Testing

or, Debunking the “Expensive Structural Types” Myth Unpublished technical memo, MIT Laboratory for Computer Science, May 07, 1991

Internet URL: http://www.swiss.csail.mit.edu/~mrb/papers.html

[Bondorf 89] Anders Bondorf A Self-Applicable Partial Evaluator for Term Rewriting Systems In [Díaz & Orejas 89], pp.81–95

Internet URL:

   http://repository.readscheme.org/ftp/papers/bondorf-term-rewriting.ps.gz
   http://repository.readscheme.org/ftp/papers/bondorf-term-rewriting.pdf

[Bondorf & Danvy 90/91] Anders Bondorf and Olivier Danvy Automatic Autoprojection of Recursive Equations with Global Variables

and Abstract Data Types

Technical Report 90/4, DIKU, University of Copenhagen, Denmark, March 1990

Superseded by Science of Computer Programming, Vol. 16, 1991, pp.151–195

Internet URL:

   ftp://ftp.daimi.au.dk/pub/empl/danvy/Papers/bondorf-danvy-scp91.ps.gz

[Boole 1854,1951/58,2003] George Boole, LL.D. An Investigation of The Laws of Thought,

on which are Founded the Mathematical Theories of Logic and Probabilities London: Walton and Maberly, 1854 (original printing) New York, N.Y.: Dover Publications, Inc., 1951, 1958 (New Edition)

Internet URL: http://www.gutenberg.org/etext/15114

Reproduced as:

John Corcoran (Introduction)

The Laws of Thought

(Great Books in Philosophy, No. 35)

Prometheus Books, July 2003

[Bower 1386/1857/1901/1966,1980,2000,2003,2005] John Bower Confessio Amantis, or, Tales of the Seven Deadly Sins Written c. 1386, first published 1390

Reprinted as Confessio Amantis of John Gower Dr. Reinhold Pauli (Editor and Collator) London: Bell and Daldy, 1857

Internet URL:

      http://books.google.com/books?id=a1sLAAAAIAAJ
      http://books.google.com/books?id=zFsLAAAAIAAJ
      http://books.google.com/books?id=FFwLAAAAIAAJ
      http://books.google.com/books?id=n1wLAAAAIAAJ

In The Works of John Gower George Campbell Macaulay (Editor) Oxford: Clarendon Press, 1901

    Internet URL: http://www.gutenberg.org/etext/266

Reprinted as Confessio Amantis (“The Lover’s Confession”)

Russell A. Peck (Editor)

Holt, Rinehart, and Winston, 1966 University of Toronto Press, 1980 Kalamazoo: Medieval Institute Publications, 2000, 2003, 2005

    Internet URL: http://books.google.com/books?id=_W-HKRVsfoQC

[Brachman 85] Ronald J. Brachman “I lied about the trees”, or, Defaults and Definitions in Knowledge Representation In AI Magazine, Vol. 6, No. 3, 1985, pp.80–93 [Brandt 97]

Michael Brandt Recursive Subtyping: Axiomatizations and Computational Interpretations Masters Thesis, TOPPS D-352, DIKU, University of Copenhagen, Denmark, Aug 1997

Internet URL: ftp://ftp.diku.dk/diku/semantics/papers/D-352.ps.gz

[Brooks 75] Frederick P. Brooks The Mythical Man-Month: Essays on Software Engineering Addison-Wesley. Reading, MA., 1975

[de Bruijn 72,95] Nicolaas Govert de Bruijn Lambda Calculus Notation with Nameless Dummies,

A Tool for Automatic Formula Manipulation,

with Application to the Church-Rosser Theorem In 𝔓𝔯𝔬𝔠. 𝔄𝔨𝔞𝔡𝔢𝔪𝔦𝔢 𝔳𝔞𝔫 𝔚𝔢𝔱𝔢𝔫𝔰𝔠𝔥𝔞𝔭𝔭𝔢𝔫, Series A: Math. Sciences, Vol. 75, 1972, pp.381–392 In Indagationes Mathematicæ, Vol. 34, 1972, pp.381–392 In Selected Papers on Automath, R.P.Nederpelt, J.H.Geuvers and R.C.deVrijer (Editors)

Studies in Logic and The Foundations of Mathematics, Vol. 133 North Holland, 1995

[Bulmer 67/79] M. G. Bulmer Principles of Statistics Dover Publications, Inc., 1967 & 1979

[Burstall & Darlington 77] Rod M. Burstall and John Darlington A Transformation System for Developing Recusrive Programs In JACM, Vol. 24, No. 1, Jan 1977, pp.44–67

[Burstall & Lampson 84] R. Burstall and B. Lampson A Kernel Language for Modules and Abstract Data Types DEC SRC TR No. 1, Digital Equipment Corp., Systems Research Center, Sep 1, 1984

Internet URL(s):

   http://www.hpl.hp.com/techreports/Compaq-DEC/SRC-RR-1.pdf
   http://research.microsoft.com/users/blampson/35-KernelModules/
                                                35-KernelModulesAbstract.htm

In Proc. International Symposium on Semantics of Data Types (Invited Paper) Giles Kahn, David B. MacQueen and Gordon Plotkin (Editors) Lecture Notes in Computer Science, Vol. 173, pp.1–50 Springer-Verlag, 1984 [Bulyonkov 88]

Mikhail A. Bulyonkov A Theoretical Approach to Polyvariant Mixed Computation In [Bjørner, Ershov & Jones 88], pp.61–64

[Calder, Feller & Eustace 99] Brad Calder, Peter Feller and Alan Eustace Value Profiling and Optimization In Journal of Instruction Level Parallelism, Mar 1999

Internet URL: http://www.cse.ucsd.edu/~calder/papers/JILP-99-VP.pdf

[Camus 42,85/55/75,2000,2005] Albert Camus ⟨⟨Le mythe de Sisyphe: essai sur l’absurde^{⟩⟩} Paris: Gallimard, c.1942, 1985

The Myth of Sisyphus and Other Essays Knopf and Random Hall, 1955

The Myth of Sisyphus: Translation of ^{⟨⟨}Le mythe de Sisyphe^{⟩⟩}, and other essays Translated by Justin O’Brien Penguin Books, 1975, 2000, Penguin Books, Ltd., 2005

[Cardelli 88] Luca Cardelli Structural Subtyping and the Notion of Power Type In Proc. ACM POPL ’88, Jan 1988, pp.70–79

Internet URL: http://lucacardelli.name/Papers/StructuralSubtyping.pdf

[Cardelli 89/91,93] Luca Cardelli Typeful Programming DEC SRC TR No. 45, Digital Equipment Corp., Systems Research Center, May 24, 1989

Revised: Jan 1, 1993

Internet URL(s): http://www.hpl.hp.com/techreports/Compaq-DEC/SRC-RR-45.pdf
                 http://lucacardelli.name/Papers/TypefulProg.pdf

In Erich J. Neuhold and Manfred Paul (Editors):

Formal Description of Programming Concepts IX, February 1989, pp.431–507 (IFIP State-of-the-Art Reports) Springer-Verlag, 1991 [Cardelli & Wegner 85]

Luca Cardelli and Peter Wegner On Understanding Types, Data Abstraction, and Polymorphism In ACM Computing Surveys, Vol. 17, No. 4, Dec 1985, pp.471–522

Internet URL: http://lucacardelli.name/Papers/OnUnderstanding.pdf

[Chambers 92] Craig Chambers The Design and Implementation of the Self Compiler,

an Optimizing Compiler for Object-Oriented Programming Languages Ph.D. Dissertation, STAN-CS-92-1420, Stanford University, March 13, 1992

Internet URL: http://research.sun.com/self/papers/craig-thesis.html

[Chambers & Ungar 89] Craig Chambers and David Ungar Customization: Optimizing Compiler Technology for Self,

a Dynamically-Typed Object-Oriented Programming Language

In Proc. ACM PLDI ’89, Jun 1989, pp.146–160

[Chambers & Ungar 90] Craig Chambers and David Ungar Iterative Type Analysis and Extended Message Splitting:

Optimizing Dynamically-Typed Object-Oriented Programs In Proc. ACM PLDI ’90, Jun 1990, pp.150–164

[Chambers & Ungar 91] Craig Chambers and David Ungar Making Pure Object-Oriented Languages Practical In Proc. ACM OOPSLA ’91, May 1991, pp.1–15

[Chambers, Ungar & Lee 89] Craig Chambers, David Ungar and Elgin Lee An Efficient Implementation of Self,

a Dynamically-Typed Object-Oriented Language Based on Prototypes In Proc. ACM OOPSLA ’89, Oct 1989, pp.49–70

[Christensen & Glück 2004] Niels H. Christensen and Robert Glück Offline Partial Evaluation can be as Accurate as Online Partial Evaluation In ACM TOPLAS, Vol. 26, No. 1, Jan 2004, pp.191–200 [Christensen, Glück & Laursen 2000]

Niels H. Christensen, Robert Glück and Søren Laursen Binding-Time Analysis in Partial Evaluation: One Size Does Not Fit All In Proc. Perspectives of System Informatics (PSI ’99)

Third International Andrei Ershov Memorial Conference D. Bjørner, M. Broy and A. V. Zamulin (Editors) Lecture Notes in Computer Science, Vol. 1755, pp.80–92 Springer-Verlag, 2000

Internet URL: ftp://ftp.diku.dk/diku/semantics/papers/D-415.ps.gz

[Church 36,65] Alonzo Church “An unsolvable problem of elementary number theory” In American Journal of Mathematics, Vol. 58, No.2, Apr 1936, pp.345–363

Reproduced in [Davis 65].

[Church 51] Alonzo Church The Calculi of Lambda-Conversion In Annals of Mathematical Studies, Vol. 6 Princeton University Press, 1951

[Clinger & Rees 91] William Clinger and Jonathan Rees (Editors) Revised^{4} Report on the Algorithmic Language Scheme MIT/AI/Memo 848b, MIT Aritifical Intelligence Laboratory, Nov 1991 Also as: ACM Lisp Pointers, Vol. IV, No. 3, Jul 1991, pp.1–55

[Consel 93p] Charles Consel Polyvariant Binding-Time Analysis for Higher-Order, Applicative Languages In Proc. ACM SIGPLAN PEPM ’93, Copenhagen, Denmark, June 1993, pp.145–154

Internet URL: http://www.irisa.fr/lande/consel/papers/bta.ps.gz

[Consel 93s] Charles Consel A Tour of Schism: A Partial Evaluation System for Higher-Order Applicative Languages In Proc. ACM SIGPLAN PEPM ’93, Copenhagen, Denmark, June 1993, pp.66–77

Internet URL: http://www.irisa.fr/lande/consel/papers/tour-schism.ps.gz

[Consel 96/97] Charles Consel Program Adaptation based on Program Transformation In Proc. ACM Workshop on Strategic Directions in Computing Research, 1996 Also as: ACM Computing Surveys, Vol. 28, No. 4es, Dec 1996, Article No. 164 Also as: ACM SIGPLAN Notices, Vol. 32, No. 1, Jan 1997, pp.69–72

Internet URL: http://www.irisa.fr/lande/consel/papers/acm-pp.ps.gz

[Consel & Danvy 93] Charles Consel and Olivier Danvy Tutorial Notes on Partial Evaluation, a.k.a. Partial Evaluation: Principles and Perspectives In Proc. ACM POPL ’93, Jan 1993, pp.493–501

Internet URL: http://www.irisa.fr/lande/consel/papers/pe-tutorial.ps.gz
       a.k.a.: http://www.daimi.au.dk/~tdpe/doc/consel-danvy-tutorial-PE.ps

[Consel, Hornof, Noyé, Noël, Volanschi 96] Charles Consel, Luke Hornof, François Noël, Jacques Noyé, Nicolae Volanschi A Uniform Approach for Compile-Time and Run-Time Specialization INRIA Research Report 2775, January 1996

Revised version:

In [Danvy, Glück & Thiemmann 96], pp.54–72

      Internet URL: http://www.irisa.fr/lande/consel/papers/tempo-tr96.ps.gz

[Consel & Khoo 91/93] Charles Consel and Siau Cheng Khoo Parameterized Partial Evaluation In Proc. ACM PLDI ’91, Jun 1991, pp.92–106 Also as: ACM SIGPLAN Notices, Vol. 26, No. 6, Jun 1991, pp.92–106

   Internet URL: http://www.comp.nus.edu.sg/~khoosc/papers/PLDI91.pdf

As condensed from the expanded version:

In ACM TOPLAS, Vol. 15, No. 3, Jul 1993, pp.463–493

   Internet URL: http://www.comp.nus.edu.sg/~khoosc/papers/ppe-khoo.pdf

[Consel & Khoo 92/95] Charles Consel and Siau Cheng Khoo On-line & Off-line Partial Evaluation: Semantic Specifications and Correctness Proofs Tech. Rep. YALE/DCS/TR912, Yale University Department of Computer Sci., Jun 1992 In Journal of Functional Programming, Vol. 5, No. 4, Oct 1995, pp.461–500

Cambridge University Press, 1995

   Internet URL: http://www.irisa.fr/lande/consel/papers/sem-fope.ps.gz

[Cooper 90] Gregory F. Cooper Bayesian Belief-Network Inference Using Recursive Decomposition Report No. KSL 90-05, Knowledge Systems Lab., Medical Comp. Sci., Feb 8, 1990 Stanford University, Stanford, California

[Cousot & Cousot 77] Patrick Cousot and Radhia Cousot “Abstract interpretation: a unified lattice model for static analysis of programs

by construction or approximation of fixpoints”

In Proc. ACM POPL ’77, Jan 1977, pp.238–252

Internet URL: http://www.di.ens.fr/~cousot/COUSOTpapers/POPL77.shtml

[Cousot & Cousot 79] Patrick Cousot and Radhia Cousot Constructive Versions of Tarski’s Fixed Point Theorems In Pacific Journal of Mathematics, Vol. 82, No. 1, May 1979, pp.43–57

Internet URL: http://www.di.ens.fr/~cousot/COUSOTpapers/Tarski-79.shtml

[Curry 42] Haskell B. Curry The Combinatory Foundations of Mathematical Logic In Journal of Symbolic Logic, Vol. 7, No. 2, Jun 1942, pp.49–64

[Curry & Feys 58/68] Haskell B. Curry and Robert Feys Combinatory Logic I. Studies in Logic and the Foundations of Mathematics North-Holland Publising Co., Amsterdam, 1958, 1968 (2nd edition)

[Danvy 96] Olivier Danvy Type-Directed Partial Evaluation In Proc. ACM POPL ’96, Jan 1996, pp.242–257

Internet URL: http://repository.readscheme.org/ftp/papers/brics/PB-494.ps.gz
              http://repository.readscheme.org/ftp/papers/brics/PB-494.pdf

[Danvy 96p] Olivier Danvy Pragmatics of Type-Directed Partial Evaluation In [Danvy, Glück & Thiemmann 96], pp.73–94

Internet URL: http://www.brics.dk/RS/96/15/

[Danvy 98] Olivier Danvy Online Type-Directed Partial Evaluation In Proc. Symposium on Functional and Logic Programming (FLOPS ’98), pp.271–295

Internet URL: http://www.brics.dk/RS/97/53/

[Danvy, Glück & Thiemmann 96] Olivier Danvy, Robert Glück and Peter Thiemann (Editors) Partial Evaluation— Selected Papers from the International Seminar

(Dagstuhl Castle, Germany, Feb 12–16, 1996)

Lecture Notes in Computer Science, Vol. 1110 Springer-Verlag, 1996

[Danvy, Malmkjær & Palsberg 94] Olivier Danvy, Karoline Malmkjær and Jens Palsberg The Essence of Eta-Expansion in Partial Evaluation In Proc. ACM SIGPLAN PEPM ’94, Orlando, Florida, USA, June 25, 1994, pp.11–20 Special issue of Lisp and Symbolic Computation, Vol. 8, No. 3, Sep 1995, pp.209–227

Kluwer Academic Publishers, 1995

Internet URL:

   ftp://ftp.dina.kvl.dk/pub/Staff/Peter.Sestoft/pepm94-proceedings/danvy.ps.gz

[Danvy, Malmkjær & Palsberg 95/96] Olivier Danvy, Karoline Malmkjær and Jens Palsberg Eta-Expansion Does The Trick BRICS RS-95-41, University of Århus, Århus, Denmark, Nov 1995

Superseded in: ACM TOPLAS, Vol. 18, No. 6, Nov 1996, pp.730–751

               Internet URL: http://www.brics.dk/RS/96/17/

[Danvy & Vestergaard 96] Olivier Danvy and René Vestergaard Semantics-Based Compiling: A Case Study in Type-Directed Partial Evaluation BRICS RS-96-13, University of Århus, Århus, Denmark, May 1996 In Proc. Programming Languages: Implementations, Logics, and Programs (PLILP ’96)

Herbert Kuchen and S. Doaitse Swierstra (Editors)

Lectures Notes in Computer Science, Vol. 1140, pp.182–197 Springer-Verlag, 1996

        Internet URL: http://www.brics.dk/RS/96/13/

[Davey & Priestley 2002] B. A. Davey and H. A. Priestley Introduction to Lattices and Order [2nd Edition] Cambridge University Press, 2002 [Davis 65]

Martin Davis (Editor) The Undecidable: Basic Papers on Undecidable Propositions, Unsolvable Problems

and Computable Functions

Dover Publications, Inc., 1965

[Dedekind 1888,1901/1963,1996/1999] Richard Dedekind

𝔚𝔞𝔰 𝔰𝔦𝔫𝔡 𝔲𝔫𝔡 𝔴𝔞𝔰 𝔰𝔬𝔩𝔩𝔢𝔫 𝔡𝔦𝔢 ℨ𝔞𝔥𝔩𝔢𝔫?

Translated as:

The Nature and Meaning of Numbers

In Wooster Woodruf Beman (Editor and Translator): Essays on the Theory of Numbers Dover Publications, Inc., 1963 (1901)

       Internet URL: http://www.gutenberg.org/etext/21016

Translated as:

What are numbers and what should they be?

In William B. Ewald (Editor):

From Kant to Hilbert: A Source Book in the Foundations of Mathematics (2 Vols) Oxford University Press, 1996 (hardcover) & 1999 (softcover), pp.787–832

[Dedekind 1890,1967/1977] Richard Dedekind Letter to Keferstein (1890) In [van Heijenoort 67/77], pp.98–103

[DeFalco 92] Tom DeFalco (Writer and Editor-in-Chief) Fantastic Four, Vol. 1, Issues #362–365, Mar 1992 to Jun 1992 Marvel Publications, Inc., 1992

Internet URL: http://www.marvunapp.com/Appendix/inniverse.htm

[Deming 50] William Edwards Deming Some Theory of Sampling Dover Publications, Inc., 1950

[Dershowitz 83] Nachum Dershowitz The Evolution of Programs (Progress in Computer Science, No. 5) Birkhäuser Boston, Inc., 1983 [Descartes 1637/1978,1983]

René Descartes ⟨⟨Discours de la méthode pour bien conduire sa raison, et chercher la verité dans les sciences^{⟩⟩} Published 1637

Translated into English as:

Discourse on the Method of Rightly Conducting the Reason, and Searching for Truth in the Sciences

    Internet URL: http://www.gutenberg.org/etext/59

In The Philosophical Works of Descartes (Volume I)

Elizabeth S. Haldane, C.H. LL.D., and G.R.T. Ross, M.A., D.Phil.

Cambridge: Cambridge University Press, 1978

Westford, Mass.: Murray Printing Company, 1983 (Vol. I)

[Díaz & Orejas 89] Josep Díaz and Fernado Orejas (Editors) Theory and Practice of Software Development (TAPSOFT ’89) Volume 2: Advanced Seminar on Foundations of Innovative Software Development II

and Colloquium on Current Issues in Programming Languages

Proc. of the Third International Joint Conference, Barcelona, Spain, Mar 13–17, 1989 Lecture Notes in Computer Science, Vol. 352 Springer-Verlag, 1989

[Dixon & Massey 57] Wilfrid J. Dixon and Frank J. Massey, Jr. Introduction-to-Statistical-Analysis [Second Edition] McGraw-Hill Book Company, Inc., 1957

[Drake 67] Alvin W. Drake Fundamentals of Applied Probability Theory (McGraw-Hill Series in Probability and Statistics) McGraw-Hill Book Company, Inc., 1967

[Driscoll & Orth 86] Graham C. Driscoll, Jr. and Donald L. Orth Compiling APL: The Yorktown APL Translator In IBM Journal of Research and Development, Vol. 30, No. 6, Nov 1986, pp.583–593

[Ellis 86] John R. Ellis Bulldog: A Compiler for VLIW Architectures ACM Doctoral Dissertation Award 1985 The MIT Press, 1986 [Ernst, Kaplan & Chambers 98]

Michael D. Ernst, Craig S. Kaplan, and Craig Chambers Predicate Dispatching: A Unified Theory of Dispatch In Proc. European Conference on Object-Oriented Programming (ECOOP ’98),

Brussels, Belgium, Jul 20–24, 1998, pp.186–211

[Ershov 77] Andrei P. Ershov On the partial computation principle In Information Processing Letters, Vol. 6, No. 2, Apr 1977, pp.38–41

Internet URL:

   http://ershov.iis.nsk.su/archive/eaindex.asp?lang=2&did=27505
   http://ershov.iis.nsk.su/archive/eaimage.asp?lang=2&did=27507&fileid=166796

[Ershov 78] Andrei P. Ershov Mixed computation in the class of recursive program schemata (To the memory of professor Lásló Kalmár) In Acta Cybernetica, Vol. 4, No. 1, 1978, pp.19–23

Internet URL: http://www.inf.u-szeged.hu/actacybernetica/prevvols/4_1/4_1.pdf

[Fibonacci c.1170–c.1240/1202/2002] Leonardo Fibonacci, a.k.a. Leonardo of Pisa, Leonardo Pisano & Leonardo Bonacci Liber Abaci Translated as:

Laurence E. Sigler (Translator and Contributor)

Fibonacci’s Liber Abaci: a translation into modern English of

Leonardo Pisano’s Book of Calculation

(Sources and Studies in the History of Mathematics and Physical Sciences)

Chapter II. § 12, Nō 26: How Many Pairs of Rabbits Are Created

by One Pair in One Year, pp.404–405

Springer-Verlag, New York, 2002 (hardcover) & 2003 (softcover)

Internet URL: http://books.google.com/books?isbn=0387407375

[Field & Harrison 88] Anthony J. Field and Peter G. Harrison Functional Programming (International Computer Science Series) Addison-Wesley Publishing Company, Inc., 1988 [Fischer 72,93]

Michael J. Fischer Lambda Calculus Schemata In Proc. ACM Conference on Proving Assertions about Programs, 1972, pp.104–109 In ACM SIGPLAN Notices, Vol. 7, No. 1 and SIGACT News , No. 14, Jan 1972

Updated as:

Lambda-Calculus Schemata

In [Kessler & Talcott 93], pp.259–288

Internet URL: http://citeseer.ist.psu.edu/fischer93lambdacalculus.html
              http://www.brics.dk/~hosc/local/LaSC-6-34-pp259-288.pdf

[Fisher 30,50,72] R. A. Fisher, Sc.D., F.R.S. (a.k.a. Sir Ronald Aylmer Fisher (1890–1962)) Inverse Probability In Proc. of the Cambridge Philosophical Society, Vol. 26, 1930, pp.528–535

Reproduced in:

W. Shewhart (Editor) with an index by J. Tukey

Contributions to Mathematical Statistics

John Wiley & Sons, 1950

Reproduced in:

J. H. Bennet (Editor)

Collected Papers of R. A. Fisher, Vol. 2, pp.428–436

University of Adelaide, 1972

Internet URL:

      http://digital.library.adelaide.edu.au/coll/special/fisher/84.pdf

[Fisher 83,98] Joseph A. Fisher Very Long Instruction Word Architectures and the ELI-512 In Proc. IEEE International Symposium on Computer Architecture (ISCA ’83)

Stockholm, Sweden, June 13–17, 1983, pp.140–150 Also as: ACM SIGARCH Computer Architecture News, Vol. 11, No. 3, Jun 1983

Reprinted in:

Gurindar S. Sohi (Editor)

25 Years of the International Symposia on Computer Architecture (selected papers) Barcelona, Spain, June 27–July 02, 1998, pp.263–273 ACM Press, 1998

See also the Retrospective, \underline{Ibid.}, pp.34–36

    Internet URL: http://www.hpl.hp.com/news/2005/jul-sep/VLIW_retrospective.pdf

[Fisher 92] Joseph A. Fisher Instruction-Level Parallelism (ILP) UVC Distinguished Lecture Series V (90 min.), Oct 21, 1992

[Fisher & O’Donnell 84] Joseph A. Fisher and John J. O’Donnell VLIW Machines: Multiprocessors We Can Actually Program Tech. Rep. YALE/DCS/TR298, Yale University Department of Computer Sci., Jan 1984 In Proc. IEEE CompCon Spring 84, San Francisco, Feb 27–Mar 1, 1984, pp.299–305

[Fisher & Rau 91,92] Joseph A. Fisher and B. Ramakrishna Rau Instruction-Level Parallel Processing In Science, Vol. 253, No. 5025, Sep 1991, pp.1233–1241

As condensed from:

Technical Report HPL-92-2, HP Software and Systems Laboratory, Jan 13, 1992

   Internet URL: http://www.hpl.hp.com/techreports/92/HPL-92-2.pdf

[Flanagan et al. 93] Cormac Flanagan, Amr Sabry, Bruce F. Duba, and Matthias Felleisen The Essence of Compiling with Continuations In Proc. ACM PLDI ’93, Jun 1993, pp.237–247

[Flanagan et al. 2003] Cormac Flanagan, Amr Sabry, Bruce F. Duba, and Matthias Felleisen The Essence of Compiling with Continuations (Retrospective) In 20 Years of the ACM/SIGPLAN Conference in

Programming Languages Design and Implementation, (1979–1999): A Selection, 2003

[Franklin 1758] Benjamin Franklin Poor Richard’s Almanack (1758 Edition)

Internet URL: http://usinfo.state.gov/usa/infousa/facts/loa/bf1758.htm

Excerpt: And again, he adviseth to Circumspection and Care, even in the smallest Matters, because sometimes a little Neglect may breed great Mischief; adding, For want of a Nail the Shoe was lost; for want of a Shoe the Horse was lost; and for want of a Horse the Rider was lost, being overtaken and slain by the Enemy, all for want of Care about a Horse-shoe Nail.

[Friedrich 1780/1999] Friedrich II, Preussen Konìnig (1712–1786), a.k.a. Frederick the Great of Prussia The Art of War Unpublished collected writings, c. 1780

Translated as:

Jay Luvaas (Editor and Translator) Frederick the Great on The Art of War Da Capo Press, Inc., 1999

[Frege 1879,1967/1977,1970,1972] Gottlob Frege

𝔅𝔢𝔤𝔯𝔦𝔣𝔣𝔰𝔰𝔠𝔥𝔯𝔦𝔣𝔱: 𝔢𝔦𝔫𝔢 𝔡𝔢𝔯 𝔞𝔯𝔦𝔱𝔥𝔪𝔢𝔱𝔦𝔰𝔠𝔥𝔢𝔫 𝔫𝔞𝔠𝔥𝔤𝔢𝔟𝔦𝔩𝔡𝔢𝔱𝔢 𝔉𝔬𝔯𝔪𝔢𝔩𝔰𝔭𝔯𝔞𝔠𝔥𝔢 𝔡𝔢𝔰 𝔯𝔢𝔦𝔫𝔢𝔫 𝔇𝔢𝔫𝔨𝔢𝔫𝔰

Halle a. S.: Louis Nebert, 1879

Translated as:

Stephen Bauer-Mengelberg (Translator)

Concept Notation: A formula language of pure thought,

modelled upon that of arithmetic, for pure thought

In [van Heijenoort 67/77], pp.1–82 Reprinted in Frege and Gödel: two fundamental texts in mathematical logic Harvard University Press, 1970

Translated as:

Conceptual Notation, A Formula Language of Pure Thought

Modelled upon the Formula Language of Arithmetic

In T. W. Bynum (Translator and Editor):

Gottlob Frege: Conceptual Notation, and related articles Oxford University Press, 1972 (Reprinted), pp.101–203

[Freund 52] John E. Freund, Ph.D. Modern Elementary Statistics Prentice-Hall, Inc., 1952

[Fuh & Mishra 88/90] You-Chin Fuh and Prateek Mishra Type Inference with Subtypes In Proc. European Symposium on Programming (ESOP ’88), Mar 1988, pp.94–114

Harald Ganzinger (Editor)

Lecture Notes in Computer Science, Vol. 330 Springer-Verlag, 1988

Reprinted in Theoretical Computer Science, Vol. 73, No. 2, Jun 1990, pp.155–175 Elsevier Science Publishers B. V. (North-Holland), 1990 [Futamura 71,99]

Yoshihiko Futamura Partial Evaluation of Computation Process— An Approach to a Compiler-Compiler In Systems. Computers. Controls, Vol. 2, No. 5, 1971, pp.45–50

Updated and revised:

In Higher-Order and Symbolic Computation, Vol. 12, No. 4, 1999, pp.381–391 Kluwer Academic Publishers, 1999

     Internet URL: http://www.brics.dk/~hosc/local/HOSC-12-4-pp381-391.pdf

[Futamura & Nogi 88] Yoshihiko Futamura and Kenroku Nogi Generalized Partial Computation In [Bjørner, Ershov & Jones 88], pp.133–151

[Futamura, Nogi & Takano 91] Yoshihiko Futamura, Kenroku Nogi and Akihiko Takano Essence of Generalized Partial Computation In Theoretical Computer Science, Vol. 90, No. 1, Nov 1991, pp.61–79

(Images of programming dedicated to the memory of Andrei P. Ershov)

Elsevier Science Publishers B. V. (North-Holland), 1991

[Ganzinger & Jones 85] Harald Ganzinger and Neil D. Jones (Editors) Programs as Data Objects Proceedings of a Workshop, Copenhagen, Denmark, October 17–19, 1985 Lecture Notes in Computer Science, Vol. 217 Springer-Verlag, 1986

[Gentzen 34,69,86] Gerhard Gentzen

𝔘𝔫𝔱𝔢𝔯𝔰𝔲𝔠𝔥𝔲𝔫𝔤𝔢𝔫 1𝔲𝔟𝔢𝔯 𝔡𝔞𝔰 𝔩𝔬𝔤𝔦𝔰𝔠𝔥𝔢 𝔖𝔠𝔥𝔩𝔦𝔢𝔢𝔫

In 𝔐𝔞𝔱𝔥𝔢𝔪𝔞𝔱𝔦𝔰𝔠𝔥𝔢 ℨ𝔢𝔦𝔱𝔰𝔠𝔥𝔯𝔦𝔣𝔱, Vol. 39, 1934–35, pp.176–210 & 405–431

Reproduced in:

Karel Berka and Lothar Kreiser (Editors)

𝔏𝔬𝔤𝔦𝔨−𝔗𝔢𝔵𝔱𝔢. 𝔎𝔬𝔪𝔪𝔢𝔫𝔱𝔦𝔢𝔯𝔱𝔢 𝔄𝔲𝔰𝔴𝔞𝔥𝔩 𝔷𝔲𝔯 𝔊𝔢𝔰𝔠𝔥𝔦𝔠𝔥𝔱𝔢 𝔡𝔢𝔯 𝔪𝔬𝔡𝔢𝔯𝔫𝔢𝔫 𝔏𝔬𝔤𝔦𝔨

Berlin: Akademie 4. Aufl. 1986, Seite 206–261

Reproduced in:

M. E. Szabo (Editor)

The Collected Papers of Gerhard Gentzen

North-Holland, 1969 [Gödel 34,65]

Kurt Gödel “On undecidable propositions of formal mathematical systems” Notes by S. C. Kleene & Barkley Rosser on lectures at the Inst. for Advanced Study, 1934 Mimeographed, Princeton, N.J., 30 pages

Reproduced in [Davis 65].

[Gomard 90] Carsten K. Gomard Partial Type Inference for Untyped Functional Programs (extended abstract) In Proc. ACM Conference on Lisp and Functional Programming, Jun 1990, pp.282–287

[Gomard & Jones 91] Carsten K. Gomard and Neil D. Jones A Partial Evaluator for the Untyped Lambda Calculus In Journal of Functional Programming, Vol. 1, No. 1, Jan 1991, pp.21–69

Cambridge University Press, 1991

   Internet URL: ftp://ftp.diku.dk/diku/semantics/papers/D-15.ps.gz

[Gonick & Smith 93] Larry Gonick and Woollcott Smith The Cartoon Guide to Statistics 1st HarperPerrenial edition, Harper Collins Publishers, Inc., 1993

[Goossens, Mittelback & Samarin 94] Michel Goossens, Frank Mittelback and Alexander Samarin The L^{A}TEX Companion Addison-Wesley Company, Inc., 1994

[Gordon 79] Michael J. C. Gordon The Denotational Description of Programming Languages (An Introduction) Springer-Verlag, 1979

[Gottlieb, Sackler & Tristan 78] Carl Gottlieb, Howard Sackler and Dorothy Tristan (uncredited) Jaws 2 Directed by: Jeannot Szwarc Produced by: David Brown and Richard D. Zanuck Music by: John Williams Distributor: Universal Pictures Release date: June 16, 1978 Running time: 116 minutes [Guibas & Wyatt 78]

Leonidas J. Guibas and Douglas K. Wyatt Compilation and Delayed Evaluation in APL In Proc. ACM POPL ’78, Aug 1978, pp.1–8

[Hansen 74] Gilbert Joseph Hansen Adaptive Systems for the Dynamic Run-Time Optimization of Programs Ph.D. Dissertation, Carnegie-Mellon University, March 1974

[Hansen & Wulf 76] G. Hansen and W. A. Wulf Adaptive Code Optimization In Program Optimization, Infotech State of the Art Report, 1976

[Hanson 91] Chris Hanson MIT Scheme Reference Manual MIT AI-TR-1281, MIT Artificial Intelligence Laboratory, Nov 1991

Internet URL: http://citeseer.ist.psu.edu/article/hanson95mit.html

[Haraldsson 78] Anders Haraldsson A Partial Evaluator, and Its Use for Compiling Iterative Statements in Lisp In Proc. ACM POPL ’78, Aug 1978, pp.195–202

[Harper & Mitchell 87] Robert Harper and Kevin Mitchell Introduction to Standard ML Computer Science Department, University of Edinburgh, Sep 1986

[Harrison 89] Williams Ludwell Harrison III The Interprocedural Analysis and Automatic Parallelization of Scheme Programs In Lisp and Symbolic Computation, Vol. 2, Nos. 3/4, Oct 1989, pp.179–396

Kluwer Academic Publishers, October 1989

Abstract URL: http://www.brics.dk/~hosc/vol02/03-harrison.html

[Hatcliff & Danvy 94] John Hatcliff and Olivier Danvy A Generic Account of Continuation-Passing Styles In Proc. ACM POPL ’94, Jan 1994, pp.17–21 [Hatcliff & Danvy 96/97]

John Hatcliff and Olivier Danvy Thunks and the λ-Calculus BRICS RS-96-19, University of Århus, Århus, Denmark, Jun 1996 In Journal of Functional Programming, Vol. 7, No. 3, May 1997, pp.303–319

Cambridge University Press, 1997

   Internet URL: http://www.brics.dk/RS/96/19/

As condensed from:

Thunks and the λ-Calculus (Extended Version)

BRICS RS-97-7, University of Århus, Århus, Denmark, Mar 1997

   Internet URL: http://www.brics.dk/RS/97/7/

[Hatcliff & Glück 96] John Hatcliff and Robert Glück Reasoning about Hierarchies of Online Program Specialization Systems In [Danvy, Glück & Thiemmann 96], pp.161–182

Internet URL: ftp://ftp.diku.dk/diku/semantics/papers/D-269.ps.gz

[Hatcliff, Torben & Thiemmann 98/99] John Hatcliff, Torben Æ. Mogensen and Peter Thiemann (Editors) Partial Evaluation: Practice and Theory (DIKU 1998 International Summer School, June 29–July 10, 1998) Lecture Notes in Computer Science, Vol. 1706 Springer-Verlag, 1999

[Hayes-Roth 88] Barbara Hayes-Roth Making Intelligent Systems Adaptive Stanford University Tech. Report STAN-CS-88-1126 and/or KSL-88-64, Oct 1988

[van Heijenoort 67/77] Jean van Heijenoort (Editor) From Frege to Gödel: A Source Book in Mathematical Logic, 1879–1931 (Source Books in the History of the Science) Harvard University Press, 1967, 1977 (3rd printing)

[Heintze 92] Nevin Heintze Set Based Program Analysis Ph.D. Dissertation, CMU-CS-92-201, Carnegie-Mellon University, 1992

Internet URL: http://citeseer.ist.psu.edu/heintze92set.html

[Heintze & Jaffar 91] Nevin Heintze and Joxan Jaffar Set-Based Program Analysis (Extended Abstract)

Internet URL: http://www.cs.cmu.edu/afs/cs/user/nch/ftp/sba.ps.Z

[Henderson 80] Peter Henderson Functional Programming: Application and Implementation (Prentice-Hall International Series in Computer Science) C. A. R. Hoare (Series Editor) Prentice-Hall International, Inc., 1980

[Henglein 91] Fritz Henglein Efficient Type Inference for Higher-Order Binding Time Analysis In [Hughes 91], pp.448-472

Internet URL: http://repository.readscheme.org/ftp/papers/henglein-bta91.pdf
              http://repository.readscheme.org/ftp/papers/henglein-bta91.ps.gz

[Hennessy & Patterson 2005/2007] John L. Hennessy and David A. Patterson Computer Organization and Design: The Hardware/Software Interface [Third Edition] Elsevier/Morgan Kaufman Publishers, Inc., 2005, (Revised Printing) 2007

[Hoare 69] C. A. R. Hoare An Axiomatic Basis for Computer Programming In CACM, Vol. 12, No. 10, Oct 1969, pp.576–580,583

[Hölzle 94] Urs Hölzle Adaptive Optimization for Self: Reconciling High Performance

with Exploratory Programming

Ph.D. Dissertation, STAN-CS-TR-94-1520, Stanford University Dept. of Comp. Sci., 1994

[Hölzle, Chambers & Ungar 91] Urs Hölzle, Craig Chambers and David Ungar Optimizing Dynamically-Typed Object-Oriented Languages

with Polymorphic Inline Caches In Proc. European Conference on Object Oriented Programming (ECOOP ’91), Geneva

Lecture Notes in Computer Science, Vol. 512, pp.21–38 Springer-Verlag, 1991 [Hopcroft 71]

John E. Hopcroft “An n lg n algorithm for minimizing states in a finite automaton” In Theory of Machines and Computations, P. Kohavi and A. Paz (Editors), pp.189–196

Academic Press, 1971

[Hopcroft & Ullman 79] John E. Hopcroft and Jeffrey D. Ullman Introduction to Automata Theory, Languages, and Computation Addison-Wesley, 1979

[Houston & Traven 48] John Houston and B. Traven The Treasure of the Sierra Madre Directed by: John Houston Produced by: Warner Bros. Written by: B. Traven (novel), John Houston (screenplay) Distributor: Warner Bros. Release date: January 6, 1948 Running time: 126 min. Budget: $3,800,000 (estimated)

Adapted from the novel: [Traven 27,35]

Original Quote URL: http://en.wikipedia.org/wiki/Stinking_badges
Video Stream URL: http://www.youtube.com/watch?v=HaxURLFn6jU
Variations  URL: http://www.darryl.com/badges

[Hudak, Wadler, et al. 88] Paul Hudak, Philip Wadler (Editors) Report on the Programming Language Haskell: Draft Proposed Standard Research Rep. YALEU/DCS/RR-666, Yale University Dept. of Computer Sci., Dec 1988

[Huff & Geis 54/82/93] Darrell Huff, Illustrated by Irving Geis How to Lie with Statistics Darrel Huff and Irving Geis, 1954 (renewed 1982) W. W. Norton & Company, Inc., 1993 (Norton paperback reissue)

[Hughes 91] John Hughes (Editor) Functional Programming Languages and Computer Architecture (FPCA ’91) (Proceedings of the 5th ACM Conference, Cambridge, MA, USA, August 26–30, 1991) Lecture Notes in Computer Science, Vol. 523 Springer-Verlag, 1991 [IEEE 91]

IEEE Scheme Standardization Committee IEEE Standard for the Scheme Programming Language IEEE Std 1178-1990, May 1991

[Ingerman et al. 61] Peter Z. Ingerman Thunks, a way of compiling procedure statements

with some comments on procedure declarations In CACM, Vol. 4, No. 1, Jan 1961, pp.55–58

From a homonymous unpublished manuscript written by Peter Ingerman, Edgar Irons, Kirk Sattley, and Wallace Feurzeig; assisted by M. Lind, Herbert Kanner, and Robert Floyd, 1960

[Jain 91] Raj Jain The Art of Computer Systems Analysis: Techniques for Experimental Design,

Measurement, Simulation, and Modeling

John Wiley & Sons, Inc., 1991

[Johnston 79] Ronald L. Johnston The Dynamic Incremental Compiler of APL\3000 In Proc. ACM APL ’79, APL Quote Quad, Vol. 9, No. 4–Part 1, Jun 1979

[Jones 90] Neil D. Jones Partial Evaluation, Self-Application and Types In Proc. Int. Colloq. on Automata, Languages and Programming (ICALP ’90), pp.639– 659

(17th International Colloquium, Warwick University, England, July 16–20, 1990) Michael S. Paterson (Editor) Lecture Notes in Computer Science, Vol. 443 Springer-Verlag, July 1990

[Jones 96] Neil D. Jones An Introduction to Partial Evaluation In ACM Computing Surveys, Vol. 28, No. 3, Sep 1996, pp.480–503

Internet URL: ftp://ftp.diku.dk/diku/semantics/papers/D-294.ps.gz

[Jones 97] Neil D. Jones Combining Abstract Interpretation and Partial Evaluation (brief overview) In P. Van Hentenryck (Editor):

Static Analysis

Lecture Notes in Computer Science, Vol. 1302, pp.396–405 Springer-Verlag. 1997

Internet URL: ftp://ftp.diku.dk/diku/semantics/papers/D-333.ps.gz

[Jones, Gomard, Bondorf, Danvy & Mogensen 90]

Neil D. Jones, Carsten K. Gomard, Anders Bondorf, Olivier Danvy

and Torben Æ. Mogensen

A Self-applicable Partial Evaluator for the Lambda Calculus (extended abstract) In Proc. IEEE Int. Conf. on Computer Languages (ICCL ’90), Mar 1990, pp.49–58

[Jones, Gomard & Sestoft 93] Neil D. Jones, Carsten K. Gomard and Peter Sestoft Partial Evaluation and Automatic Program Generation (Prentice Hall International Series in Computer Science) Prentice Hall International (UK) Ltd, 1993

For full text download, plus support materials:

   http://www.dina.kvl.dk/~sestoft/pebook/
   http://www.dina.kvl.dk/~sestoft/pebook/jonesgomardsestoft.ps
   http://www.dina.kvl.dk/~sestoft/pebook/jonesgomardsestoft-letter.pdf
   http://www.dina.kvl.dk/~sestoft/pebook/jonesgomardsestoft-a4.pdf

To browse & search the full book content:

   http://books.google.com/books?id=7rPPScYo8w8C

[Jones & Schmidt 80] Neil D. Jones and David A. Schmidt Compiler Generation from Denotational Semantics In Neil D. Jones (Editor):

Semantics-Directed Compiler Generation, Århus, Denmark Lecture Notes in Computer Science, Vol. 94, pp.70–93 Springer-Verlag, 1980

[Jones, Sestoft & Søndergaard 89] Neil D. Jones, Peter Sestoft and Harald Søndergaard Mix: A Self-Applicable Partial Evaluator for Experiments in Compiler Generation In Lisp and Symbolic Computation, Vol. 2, Nos. 1, Feb 1989, pp.9–50

Kluwer Academic Publishers, October 1989

Internet URL: http://www.brics.dk/~hosc/local/LaSC-2-1-pp9-50.pdf

[Jørgensen 98] Jesper Jørgensen. Similix: A Self-Applicable Partial Evaluator for Scheme In [Hatcliff, Torben & Thiemmann 98/99], pp.83–107

Internet URL: http://www.dina.dk/~jesper/PEsummerschool/Similix/similix.html

[Kahn 60,69,78,2003,2007] Herman Kahn On Thermonuclear War Princeton Univeristy Press, 1960 Free Press, New York, NY, 1969 Greenwood Press, 1978 Textbook Publishers, 2003 Transaction Publishers, 2007 (contributor: Evan Jones)

[Kearns & Vazirani 94] Michael J. Kearns and Umesh V. Vazirani An Introduction to Computational Learning Theory The MIT Press, 1994

[Kelsey, Clinger & Rees 98] Richard Kelsey, William Clinger and Jonathan Rees (Editors) Revised^{5} Report on the Algorithmic Language Scheme In ACM SIGPLAN Notices, Vol. 33, No. 9, Sep 1998, pp.26–76 In Higher-Order and Symbolic Computation, Vol. 11, No. 1, Aug 1998, pp.7–105

Internet URL: http://www.brics.dk/~hosc/local/HOSC-11-1-pp7-105.pdf

[Keppel 91] David Keppel A Portable Interface for On-The-Fly Instruction Space Modification In Proc. ACM ASPLOS ’91, Apr 1991, pp.86–95

[Keppel, Eggers & Henry 91] David Keppel, Susan J. Eggers, and Robert R. Henry A Case for Runtime Code Generation TR 91-11-04, University of Washington, Dept. of Comp. Sci. & Eng., 1991

[Kesselman 91] Carl Kesselman Tools and Techniques for Performance Measurement and Performance Improvement

in Parallel Programs

Ph.D. Dissertation, UCLA CSD-910015, Univ. California at Los Angeles, June 1991 [Kessler & Talcott 93]

Robert R. Kessler and Carolyn Talcott (Editors) Lisp and Symbolic Computation,, Vol. 6, Nos. 3–4, 1993 (Special issue on continuations— part I) Kluwer Academic Publishers, 1993

Internet URL: http://www.brics.dk/~hosc/vol06/contents.html

[Kiczales, des Rivières and Bobrow 91] Gregor Kiczales, Jim des Rivières and Daniel G. Bobrow

The Art of the Metaobject Protocol

The MIT Press, 1991

[Kleene 50] Stephen Cole Kleene Introduction to Metamathematics D. Van Nostrand Company, Inc., 1950

[Knuth 74] Donald E. Knuth Structured Programming with go to Statements In ACM Computing Surveys, Vol. 6, No. 4, Dec 1974, pp.261–301

“We should forget about small efficiencies, say about 97% of the time: premature optimization is the root of all evil.” [p.268]

[Knuth 86] Donald E. Knuth (Illustrations by Duane Bibby) The TEXbook Addison Wesley Publishing Company, 1986

[Krall 94] Andreas Krall Improving Semi-Static Branch Prediction by Code Replication In Proc. ACM PLDI ’94, Jun 1994, pp.97–106 [Kubrick 64]

Stanley Kubrick (Director) Dr. Strangelove, or: How I Learned to Stop Worrying and Love the Bomb Directed by Stanley Kubrick Produced by Stanley Kubrick Novel: Peter George Screenplay: Stanley Kubrick, Peter George, Terry Southern,

Peter Sellers (uncredited), James B. Harris (uncredited)

Distributed by Columbia Pictures Release date: January 29, 1964 Running time: 94 min. Country: United Kingdom & United States Budget: $1,800,000

[Kyrimis & Alonso 88] Kriton Kyrimis and Rafael Alonso An Experimental Comparison of Initial Placement vs. Process Migration

for Load Balancing Strategies

Princeton University CS-TR-199-88, Dec 1988

[Lampson & Burstall 88] B. Lampson and R. Burstall Pebble, a Kernel Language for Modules and Abstract Data Types In Information and Computation, Vol. 76, Nos. 2/3, Feb/Mar 1988, pp.278–346

Internet URL:

   http://research.microsoft.com/users/blampson/39-Pebble/39-PebbleAbstract.htm

[Lamport 86] Leslie Lamport (Illustrations by Duane Bibby) L^{A}TEX: A Document Preparation System Addison-Wesley Publishing Company, Inc., 1986

[Landin 64] Peter J. Landin The Mechanical Evaluation of Expressions In The Computer Journal, Vol. 6, No. 4, 1964, pp.308–320

See also [Henderson 80] & [Field & Harrison 88].

[Langley 70] Dr. Russell Langley Practical Statistics Simply Explained [Revised Edition] Dover Publications, Inc., 1970 [Larus 90]

James R. Larus Abstract Execution: A Technique for Efficiently Tracing Programs In Software— Practice and Experience, Vol. 20, No. 12, Dec 1990, pp.1241–1258

[Launchbury 91] John Launchbury Projection Factorizations in Partial Evaluation Distinguished Dissertations in Computer Science, Cambridge University Press, 1991

[Lee 89] Spike Lee (Writer, Producer and Director) Do the Right Thing Distributor: Universal Pictures Releaseed: May, 1989 (première at Cannes),

June 30, 1989 (public release)

Running time: 125 minutes Budget: $6,500,000

[von Leibniz 1686] Gottfried Wilhelm Freiherr von Leibniz ⟨⟨Discours de Métaphysique^{⟩⟩}

Translated in:

Dr. George R. Montgomery (Translator) Paul Janet (Introduction)

Discourse on Metaphysics, Correspondence with Arnauld and Monadology The Open Court Publishing Company, 1902, 1908, 1916 Kessinger Publishing, 2003

    Internet URL: http://books.google.com/books?q=isbn:0766136604

Reprinted as:

George R. Montgomery (Translator)

Discourse on Metaphysics and the Monadology Prometheus Books, 1992

Internet URL(s):

   http://www.4literature.net/Gottfried_Wil_Leibniz/Discourse_on_Metaphysics/
   http://www.earlymoderntexts.com/pdf/leibdm.pdf

[Lisberger & MacBird 82] Steven Lisberger and Bonnie MacBird Tron Directed by: Steven Lisberger Produced by: Donald Kushner Written by: Steven Lisberger & Bonnie MacBird (story), Steven Lisberger (screenplay) Distributed by: Buena Vista Pictures Release date: July 9, 1982 (USA) Running time: 96 min. Budget: $17,000,000 (estimated) Gross revenue: $33,000,000 (USA)

[Lombardi 67] Lionello Lombardi Incremental Computation In Advances in Computers, Vol. 8, 1967, pp.247–333

[Lombardi & Raphael 64/66] Lionello A. Lombardi and Bertram Raphael Lisp as the Language for an Incremental Computer In Edmund C. Berkeley and Daniel G. Bobrow (Editors):

The Programming Language Lisp: Its Operation and Applications, pp.204–219 Information International, Inc., March, 1964 The MIT Press, Apr 1966

Internet URL:

   http://community.computerhistory.org/scc/projects/LISP/book/
          III_LispBook_Apr66.pdf

[Lumer 91] Erik Lumer Ultradiffusion Algorithm for Resource Management In Distributed Systems Xerox PARC SSL-91-103, System Sciences Lab., Xerox Palo Alto Research Center, 1991

[McCarthy 60] John McCarthy Recursive Functions of Symbolic Expressions and Their Computation by Machine, Part I In CACM, Vol. 3, No. 4, 1960, pp.184–195

[McCarthy et al. 62]

John McCarthy, Paul W. Abrahams, Daniel J. Edwards, Timothy P. Hart

and Michael I. Levin

Lisp 1.5 Programmer’s Manual (Second Edition) The MIT Press, 1962 Internet URL:

   http://community.computerhistory.org/scc/projects/LISP/book/
          LISP%201.5%20Programmers%20Manual.pdf

[Malmkjær 89] Karoline Malmkjær Program and Data Specialization Principles, Applications, and Self-Application Masters Thesis (“Speciale”), DIKU, University of Copenhagen, Denmark, August 1989

Internet URL:

   http://repository.readscheme.org/ftp/papers/malmkjaer-speciale.ps.gz
   http://repository.readscheme.org/ftp/papers/malmkjaer-speciale.pdf

[Malmkjær & Ørbæk 95] Karoline Malmkjær and Peter Ørbæk Polyvariant Specialisation for Higher-Order, Block-Structured Languages In Proc. ACM SIGPLAN PEPM ’95, La Jolla, California, June 21–23, 1995, pp.66–76

Internet URL: ftp://ftp.daimi.au.dk/pub/empl/poe/pepm95.ps.gz

[Mandel 64] John Mandel The Statistical Analysis of Experimental Data John Wiley & Sons, Inc., 1964

[Martel 2006] Matthieu Martel Semantics of roundoff error propagation in finite precisionn computations In Higher-Order and Symbolic Computation, Vol. 19, No. 1, 2006, pp.7–30

Kluwer Academic Publishers, 2006

   Internet URL: http://www.enseignement.polytechnique.fr/informatique/profs/
                        Matthieu.Martel/hosc04.pdf

[Martel 2005] Matthieu Martel An Overview of Semantics for the Validation of Numerical Programs In Proc. Verification, Model Checking, and Abstract Interpretation (VMCAI ’05)

Radhia Cousot (Editor)

Lecture Notes in Computer Science, Vol. 3385, pp.59–77 Springer-Verlag, 2005

Internet URL:

                 http://www.enseignement.polytechnique.fr/informatique/profs/
                        Matthieu.Martel/vmcai05.pdf

[Massalin & Pu 90] Henry Massalin and Calton Pu Fine-Grain Adaptive Scheduling using Feedback In USENIX Computing Systems, Vol. 3, No. 1, Winter 1990, pp.139–173

[Mazurkiewicz 71] Antoni W. Mazurkiewicz Proving Algorithms by Tail Functions In Information and Control, Vol. 18, No. 3, 1971, pp.220–226

(The final version of the paper was printed two years later than the original.)

[Might 2007] Matthew Might Environment Analysis of Higher-Order Languages Ph.D. Dissertation, Georgia Institute of Technology, August 2007

Internet URL: http://www.smartech.gatech.edu/bitstream/1853/16289/1/
                     might_matthew_b_200708_phd.pdf

[Might & Shivers 2006] Matthew Might and Olin Shivers Environment Analysis via ΔCFA In Proc. ACM POPL ’2006, Jan 2006, pp.127–140 Also as: ACM SIGPLAN Notices, Vol. 41, No. 1, Jan 2006, pp.127–140

Internet URL: http://matt.might.net/papers/might2006dcfa.pdf
              http://www.ccs.neu.edu/home/shivers/papers/dcfa.ps

[Might & Shivers 2007] Matthew Might and Olin Shivers Analyzing the Environment Structure of Higher-Order Languages using Frame Strings In Theoretical Computer Science, Vol. 375, Nos. 1–3, May 2007, pp.137–168

North-Holland Publishing Company, 2006

Internet URL: http://matt.might.net/papers/might2007dcfa.pdf
              http://www.ccs.neu.edu/home/shivers/papers/dcfa-tcs.ps

[Miller & Rozas 91] James S. Miller and Guillermo J. Rozas Free Variables and First-Class Environments In Lisp and Symbolic Computation: An International Journal, Vol. 3, No. 4, pp.107–144

Kluwer Academic Publishers, April 1991

Abstract URL: http://www.brics.dk/~hosc/vol04/02-miller-rozas.html

[Milner 84] Robin Milner A Proposal for Standard ML In Proc. ACM Symposium on Lisp and Functional Programming, Aug 1984, pp.184–197

[Mitchell 81] Tom M. Mitchell Generalization as Search In Bonnie Lynn Webber and Nils J. Nilson (Editors):

Readings in Artificial Intelligence, pp.517–542 Morgan Kaufman Publishers, Inc., 1981

[Mitchell 84] John C. Mitchell Coercion and Type Inference (Summary) In Proc. ACM POPL ’84, Jan 1984, pp.175–185

[Mogensen 88] Torben Ægidius Mogensen Partially Static Structures in a Self-Applicable Partial Evaluator In [Bjørner, Ershov & Jones 88], pp.325–347

[Mogensen 89] Torben Æ. Mogensen Binding Time Aspects of Partial Evaluation Ph.D. Dissertation, DIKU, University of Copenhagen, Denmark, March 1989

[Mogensen 92] Torben Æ. Mogensen Self-applicable Partial Evaluation for Pure Lambda Calculus In Proc. ACM SIGPLAN PEPM ’92, San Francisco, California, June 1992, pp.116–121

Internet URL: http://repository.readscheme.org/ftp/papers/topps/D-141.ps.gz
              http://repository.readscheme.org/ftp/papers/topps/D-141.pdf

[Mogensen 95] Torben Æ. Mogensen Self-Applicable Online Partial Evaluation of the Pure Lambda Calculus In Proc. ACM SIGPLAN PEPM ’95, La Jolla, California, June 21–23, 1995, pp.39–44

[Mogensen 96] Torben Æ. Mogensen Evolution of Partial Evaluators: Removing Inherited Limits In [Danvy, Glück & Thiemmann 96], pp.303–321 [Morris 70/93]

Lockwood Morris The Next 700 Formal Language Descriptions In [Kessler & Talcott 93], pp.249–256

(Original manuscript dated Nov 1970)

Internet URL: http://www.brics.dk/~hosc/local/LaSC-6-34-pp249-258.pdf

[Mulmuley 86] Ketan Mulmuley Full Abstraction and Semantic Equivalence ACM Doctoral Dissertation Award 1986 The MIT Press, 1986

[Mycroft 81] Alan Mycroft Abstract Interpretation and Optimising Transformations of Applicative Programs Ph.D. Dissertation, CST-15-81, Edinburgh University, 1981

[von Neumann 45,93] John von Neumann First Draft of a Report on the EDVAC Contract No. W-670-ORD-4926, University of Pennsylvania, June 30, 1945

Reproduced in:

Michael D. Godfrey (Editor, with Introduction)

IEEE Annals of the History of Computing, Vol. 15, No. 4, 1993, pp.27–75

Internet URL: http://qss.stanford.edu/~godfrey/vonNeumann/vnedvac.pdf
              http://www.virtualtravelog.net/entries/2003-08-TheFirstDraft.pdf

[von Neumann 46,81,88] John von Neumann The Principles of Large-Scale Computing Machines In IEEE Annals of the History of Comp., Vol. 3, No. 3, Jul–Sep 1981, pp.263–273 In IEEE Annals of the History of Comp., Vol. 10, No. 4, Oct–Dec 1988, pp.243–256

[Nemytykh, Pinchuk & Turchin 96] Andrei P. Nemytykh, Victoria A. Pinchuk and Valentin F. Turchin A Self-Applicable Supercompiler In [Danvy, Glück & Thiemmann 96], pp.322–337

   Internet URL: http://www.citeseer.ist.psu.edu/50162.html

[Nielson 86] Flemming Nielson Abstract Interpretation of Denotational Definitions (A survey) In Symposium on Theoretical Aspects of Computer Science (STACS ’86), Orsay, France

Lecture Notes in Computer Science, Vol. 210, pp.1–20 Springer-Verlag, 1986

[Nielson 87,88] Flemming Nielson Strictness Analysis and Denotational Abstract Interpretation (Extended Abstract) In Proc. ACM POPL ’87, Jan 1987, pp.120–131

As condensed from:

Information and Computation, Vol. 76, No. 1, Jan 1988, pp.29–92

[Nielson & Nielson 88,88] Hanne R. Nielson and Flemming Nielson Automatic Binding Time Analysis for a Typed λ-Calculus (Extended Abstract) In Proc. ACM POPL ’88, Jan 1988, pp.98–106

As condensed from:

Science of Computer Programming, Vol. 10, No. 2, Apr 1998, pp.139–176 Elsevier Science Publishers B. V. (North-Holland), 1988

[Nielson & Nielson 91,92] Hanne Riis Nielson and Flemming Nielson Bounded Fixed Point Iteration (Extended Abstract) In Proc. ACM POPL ’92, Jan 1992, pp.71–82

As condensed from:

Journal of Logic and Computation, Vol. 2, No. 4, Aug 1992, pp.441–464 Oxford University Press, 1992

As condensed from:

Tech. Rep. DAIMI PB-359, Computer Sci. Dept., Århus University, Denmark, Jul 1991

   Internet URL: http://www.daimi.au.dk/PB/359/PB-359-abstract.html
                 http://www.daimi.au.dk/PB/359/PB-359.pdf

[Nielson & Nielson 92] Flemming Nielson and Hanne Riis Nielson Two-Level Functional Languages In Cambridge Tracts in Theoretical Computer Science, No. 34

Cambridge University Press, 1992 [Nielson & Nielson 92,92]

Flemming Nielson and Hanne Riis Nielson Finiteness Conditions for Fixed Point Iteration (Extended Abstract) In Proc. ACM Conference on Lisp and Functional Programming, Jun 1992, pp.96–108 Also as: ACM Lisp Pointers, Vol. V, No. 1, Jan. 1992, p.96-108

As condensed from:

Tech. Rep. DAIMI PB-384, Computer Sci. Dept., Århus University, Denmark, Feb 1992

   Internet URL: http://www.daimi.au.dk/PB/384/PB-384-abstract.html
                 http://www.daimi.au.dk/PB/384/PB-384.ps.gz
                 http://www.daimi.au.dk/PB/384/PB-384.pdf

[Nielson & Nielson 93,93] Flemming Nielson and Hanne Riis Nielson Finiteness Conditions for Strictness Analysis (Extended Abstract) In Proc. Third International Workshop on Static Analysis (WSA ’93), Sep 1993

Patrick Cousot, Moreno Falaschi, Gilberto Filè and Antoine Rauzy (Editors) Lecture Notes in Computer Science, Vol. 724, pp.194–205 Springer-Verlag, 1993

As condensed from:

Tech. Rep. DAIMI PB-447, Computer Sci. Dept., Århus University, Denmark, Jul 1993

   Internet URL: http://www.daimi.au.dk/PB/447/PB-447-abstract.html
                 http://www.daimi.au.dk/PB/447/PB-447.pdf

[Nielson & Nielson 96] Flemming Nielson and Hanne Riis Nielson Multi-Level Languages: a Descriptive Framework Tech. Rep. DAIMI PB-510, Computer Sci. Dept., Århus University, Denmark, Dec 1996

Internet URL: http://www.daimi.au.dk/PB/510/PB-510-abstract.html
              http://www.daimi.au.dk/PB/510/PB-510.ps.gz
              http://www.daimi.au.dk/PB/510/PB-510.pdf

[Niven & Pournelle 74] Larry Niven and Jerry Pournelle The Mote in God’s Eye Simon and Schuster, 1974

[Noël, Hornof, Consel & Lawall 98] François Noël, Luke Hornof, Charles Consel and Julia L. Lawall Automatic, Template-Based Run-Time Specialization: Implementation

and Experimental Study

In Proc. IEEE Int. Conf. on Computer Languages (ICCL ’98), May 1998, pp.123–142

Internet URL: http://www.irisa.fr/lande/consel/papers/rt_bench.ps.gz

[Palsberg 98] Jens Palsberg Eta-Redexes in Partial Evaluation In [Hatcliff, Torben & Thiemmann 98/99], pp.356–366

Internet URL:

   http://repository.readscheme.org/ftp/papers/pe98-school/palsberg-pe98.ps.gz
   http://repository.readscheme.org/ftp/papers/pe98-school/palsberg-pe98.pdf

[Palsberg & Schwartzbach 92,94] Jens Palsberg and Michael I. Schwartzbach Binding Time Analysis: Abstract Interpretation versus Type Inference Tech. Rep. DAIMI PB-393, Computer Sci. Dept., Århus University, Denmark, Apr 1992

Internet URL: http://www.daimi.au.dk/PB/393/PB-393-abstract.html
              http://www.daimi.au.dk/PB/393/PB-393.ps.gz
              http://www.daimi.au.dk/PB/393/PB-393.pdf

Updated as:

Binding-time Analysis: Abstract Interpretation versus Type Inference In Proc. IEEE Int. Conf. on Computer Languages (ICCL ’94), May 1994, pp.289–298

      Internet URL: http://citeseer.ist.psu.edu/palsberg94bindingtime.html

[Patterson & Hennessy 2006] David A. Patterson and John L. Hennessy Computer Architecture: A Quantitative Approach [Fourth Edition] Elsevier/Morgan Kaufman Publishers, Inc., 2006

[Peano 1889,1967/1977] Giuseppe Peano Arithmetices principia, nova methodo exposita

Translated as:

The principles of arithmetic, presented by a new method In [van Heijenoort 67/77], pp.83–97

[Perlis 82] Alan J. Perlis, Yale University Epigrams on Programming In ACM SIGPLAN Notices, Vol. 17, No. 9, Sep 1982, pp.7–13

Internet URL: http://www.cs.yale.edu/quotes.html

[Pettis & Hansen 90] Karl Pettis and Robert C. Hansen Profile Guided Code Positioning In Proc. ACM PLDI ’90, Jun 1990, pp.16–27

[Phillips 96] John L. Phillips, Jr. How to Think about Statistics [5th Edition] W. H. Freeman and Company, 1996

[Pingala 450–200BC] pin.galá (Sanskrit grammarian) chandah.-śāstra (The Art of Prosody)

Note: Contains the basic ideas of Fibonacci numbers (called mātrāmeru, or “mountain of cadence”)

Internet URL: http://en.wikipedia.org/wiki/Fibonacci_number

[Plotkin 75] Gordon D. Plotkin Call-by-Name, Call-by-Value and the Lambda Calculus In Theoretical Computer Science, Vol. 1, No. 2, Dec 1975, pp.125–159

North-Holland Publishing Company, 1975

Internet URL: http://homepages.inf.ed.ac.uk/gdp/publications/cbn_cbv_lambda.pdf

[Plotkin 81] Gordon D. Plotkin A Structural Approach to Operational Semantics Tech. Rep. DAIMI FN-19, Computer Science Dept., Århus University, Denmark, 1981

[Post 36,65] Emil L. Post “Finite combinatory processes— formulation I” In Journal of Symbolic Logic, Vol. 1, No. 3, Sep 1936, pp.103–105

Reproduced in [Davis 65].

[Pu, Massalin & Ioannidis 88] Calton Pu, Henry Massalin and John Ioannidis The Synthesis Kernel In USENIX Computing Systems, Vol. 1, No. 1, Winter 1988, pp.11–32 [Putot, Goubault & Martel 2004]

Sylvie Putot, Eric Goubault and Matthieu Martel Static Analysis-Based Validation of Floating-Point Computations In Numerical Software with Result Verification

René Alt, Andreas Frommer, R. Baker Kearfott and Wolfram Luther (Editors)

Revised Papers from the International Seminar

(Dagstuhl Castle, Germany, January 19–24, 2003)

Lecture Notes in Computer Science, Vol. 2991, pp.306–313

Springer-Verlag, 2004

   Internet URL: http://www.enseignement.polytechnique.fr/informatique/profs/
   Internet URL:  Matthieu.Martel/DagstuhlFinal.pdf

[Python 70,73,99] John Cleese, Terry Jones, Michael Palin, Graham Chapman, Eric Idle and Terry Gilliam Monty Python’s Flying Circus, Television Series 1969–1974 (Director: Ian McNaughton) Series episode No. 22: Aired November 24, 1970; Recorded September 25, 1970 Series 2 (Sep 1970–Dec 1970), No. 9: “How to Recognise Different Parts of the Body” Production Co.: British Broadcasting Corporation (BBC) & Python (Monty) Pictures Distributor: British Broadcasting Corporation (BBC) (UK) (TV) (original airing)

Studio album: The Monty Python Matching Tie and Handkerchief by Monty Python Released 1973, Recorded September 1973, Genre: Comedy, Length 41:11, Side 1, Track 5 Label: Charisma Records/Arista Records

DVD: A&E Home Video presents Monty Python’s Flying Circus Set Four, Vol. 7 Distributors: A&E Home Video, 1999 (USA), BBC Video & BMG Video (UK) Release Date: November 16, 1999, Run Time: 01h:42m:00s, UPC: 733961700503

Video Stream URL: http://www.youtube.com/watch?v=_f_p0CgPeyA
Transcript URL: http://en.wikipedia.org/wiki/Bruces_sketch

[Python 75] Graham Chapman, John Cleese, Terry Gilliam, Eric Idle, Terry Jones and Michael Palin Monty Python and the Holy Grail Directed by: Terry Gilliam & Terry Jones Produced by: Mark Forstater, Michael White Distributor: 20th Century Fox (UK, video), Columbia Pictures (USA), EMI (UK, 1975) Releaseed: April 3, 1975 (UK) Running time: 91 min. Budget: £229,000

Video Stream URL: http://www.youtube.com/watch?v=zrzMhU_4m-g

[Quine 60] Willard Van Orman Quine Word and Object The MIT Press, 1960

[Rau & Fisher 92,93] B. Ramakrishna Rau and Joseph A. Fisher Instruction-Level Parallel Processing: History, Overview and Perspective In The Journal of Supercomputing, Vol. 2, Nos. 1–2, May 1993, pp.9–50

Kluwer Academic Publishers

Also as:

Technical Report HPL-92-132, HP Computer Systems Laboratory, Oct 26, 1992

     Internet URL: http://www.hpl.hp.com/techreports/92/HPL-92-132.pdf

[Raymond 96] Eric S. Raymond (Compiler) The New Hacker’s Dictionary [Third Edition] Compiled by Eric S. Raymond with foreword and cartoons by Guy L. Steele, Jr. The MIT Press, 1996

[Reilly 90] Matthew H. Reilly A Performance Monitor for Parallel Programs Academic Press Inc., 1990

[Reynolds 72,98] John C. Reynolds Definitional Interpreters for Higher-Order Programming Languages In Proc. ACM Annual Conference, Vol. 2, Boston, Aug 1972, pp.717–740

Reprinted with minor corrections in:

Olivier Danvy and Carolyn Talcott (Editors)

Higher-Order and Symbolic Computation, Vol. 11, No. 4, 1998, pp.363–397 Kluwer Academic Publishers, 1998

    Internet URL: http://www.brics.dk/~hosc/local/HOSC-11-4-pp363-397.pdf

See also more substantive corrections and notes in:

John C. Reynolds

Definitional Interpreters Revisited, \underline{Ibid.}, pp.355–361

    Internet URL: http://www.brics.dk/~hosc/local/HOSC-11-4-pp355-361.pdf

[Reynolds 93] John C. Reynolds, The discoveries of continuations In [Kessler & Talcott 93], pp.233–247

Internet URL: http://citeseer.ist.psu.edu/reynolds93discoveries.html
              http://www.brics.dk/~hosc/local/LaSC-6-34-pp233-248.pdf

[Rosen 68,2001] Leo Calvin Rosten The Joys of Yiddish; a relaxed lexicon of Yiddish, Hebrew and Yinglish words

often encountered in English… from the days of the Bible to those of the beatnik

New York: McGraw-Hill, 1968

Revised and updated as: The new Joys of Yiddish by Leo Rosten; revisions and commentary by Lawrence Bush;

illustrations by R. O. Blechman

New York: Crown Publishers, 2001

[Rosenkrantz & Hunt 87] D. J. Rosenkrantz and H. B. Hunt Efficient Algorithms for Automatic Construction and Compaction of Parsing Grammars In ACM TOPLAS, Vol. 9, No. 4, 1987, pp.543–566

[Roylance 87/88] Gerald L. Roylance Expressing Mathematical Subroutines Constructively MIT AI Memo No. 999, MIT Artificial Intelligence Laboratory, Nov 1987 In Proc. ACM Conference on Lisp and Functional Programming, Jul 1988, pp.8–13

[Rozas 84] Guillermo J. Rozas Liar, an ALGOL-like Compiler for Scheme S.B. Thesis, MIT Artificial Intelligence Laboratory, Jan 1984

[Rozas 92] Guillermo Juan Rozas Taming the Y Operator In Proc. ACM Conference on Lisp and Functional Programming, Jun 1992, pp.226–234

Internet URL: ftp://ftp.cs.indiana.edu/pub/scheme-repository/doc/pubs/y.ps.gz

[Rozas 93] Guillermo J. Rozas Translucent Procedures, Abstraction without Opacity Ph.D. Dissertation, MIT AI-TR-1427, MIT Artificial Intelligence Laboratory, Oct 1993

Internet URL: http://citeseer.ist.psu.edu/rozas93translucent.html

[Rozas 2007] Personal communication with Guillermo J. Rozas, March 24 & 25, 2007.

[Ruf 93] Erik Steven Ruf Topics in Online Partial Evaluation Ph.D. Dissertation, Stanford University Computer Systems Laboratory, February 1993 FUSE Memo 93-14, CSL-TR-93-563, Stanford University Comp. Systems Lab., Mar 1993

Internet URL:

   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-93-14.ps.gz
   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-93-14.pdf

[Ruf & Weise 91] Erik Ruf and Daniel Weise Using Types to Avoid Redundant Specialization FUSE MEMO 91-5, Stanford University Computer Systems Laboratory, Jun 1991

This version corrects a couple small errors in Section 2.3 of PEPM ’91, pp.321–333.

Internet URL:

   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-91-5.ps.gz
   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-91-5.pdf

Extended by [Ruf & Weise 92r].

[Ruf & Weise 92o] Erik Ruf and Daniel Weise Opportunities for Online Partial Evaluation FUSE Memo 92-7, CSL-TR-92-516, Stanford University Comp. Systems Lab., Apr 1992

Internet URL:

   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-92-7.ps.gz
   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-92-7.pdf

See also [Ruf 93, Chapter 2].

[Ruf & Weise 92p] Erik Ruf and Daniel Weise Preserving Information During Online Partial Evaluation FUSE Memo 92-8, CSL-TR-92-517, Stanford University Comp. Systems Lab., Apr 1992

Internet URL:

   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-92-8.ps.gz
   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-92-8.pdf

See also [Ruf 93, Chapters 4 & 5].

[Ruf & Weise 92r] Erik Ruf and Daniel Weise Avoiding Redundant Specialization During Partial Evaluation FUSE Memo 92-9, CSL-TR-92-518, Stanford University Comp. Systems Lab., Apr 1992

Internet URL:

   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-92-9.ps.gz
   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-92-9.pdf

Extends [Ruf & Weise 91].

See also [Ruf 93, Chapter 6].

[Ruf & Weise 92x] Erik Ruf and Daniel Weise Improving the Accuracy of Higher-Order Specialization using Control Flow Analysis FUSE MEMO 92-11, Stanford University Computer Systems Laboratory, Jun 1992 In Proc. ACM SIGPLAN PEPM ’92, San Francisco, California, June 1992, pp.67–74

Internet URL:

   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-92-11.ps.gz
   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-92-11.pdf

See also [Ruf 93, Chapter 5].

[Ruf & Weise 92,93] Erik Ruf and Daniel Weise On the Specialization of Online Program Specializers FUSE Memo 92-13, CSL-TR-92-534, Stanford University Comp. Systems Lab., Jul 1992

Internet URL:

   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-92-13.ps.gz
   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-92-13.pdf

Updated in Journal of Functional Programming, Vol. 3, No. 3, Jul 1993, pp.251–281 Cambridge University Press, 1993

           Internet URL: http://citeseer.ist.psu.edu/ruf92specialization.html

See also [Ruf 93, Chapter 7].

[Russell 08,67/77,56/88–2004] Bertrand Russell Mathematical Logic as Based on the Theory of Types In American Journal of Mathematics, Vol. 30, No. 3, Jul 1908, pp.222–262

Reprinted in [van Heijenoort 67/77], pp.152–182 Reprinted in:

Robert Charles Marsh (Editor)

Russell, Bertrand, Logic and Knowledge: Essays 1901–1950, pp.59–102 Allen and Unwin, 1956 Unwin Hyman Ltd., 1988 & 1989 Routledge, 1992, 1994, 1997, 2001, and 2004

[Rustin 72] Randall Rustin (Editor) Design and Optimization of Compilers In Proc. Courant Computer Science Symposium 5, Mar 1971

Prentice-Hall, Inc., 1972

[Ryder 74] K. D. Ryder Optimizing program placement in virtual systems In IBM Systems Journal, Vol. 13, No. 4, 1974

[Samples 91] Alan Dain Samples Profile-Driven Compilation Ph.D. Dissertation, UCB/CSD 91/627, UC Berkeley, Comp. Sci. Div. (EECS), April 1991

[Schmidt 86] David A. Schmidt Denotational Semantics: A Methodology for Language Development Allyn and Bacon, Inc., 1986

[Schönfinkel 24,67/77] Moses Schönfinkel

1 𝔘𝔟𝔢𝔯 𝔡𝔦𝔢 𝔅𝔞𝔲𝔰𝔱𝔢𝔦𝔫𝔢 𝔡𝔢𝔯 𝔪𝔞𝔱𝔥𝔢𝔪𝔞𝔱𝔦𝔰𝔠𝔥𝔢𝔫 𝔏𝔬𝔤𝔦𝔨

In 𝔐𝔞𝔱𝔥𝔢𝔪𝔞𝔱𝔦𝔰𝔠𝔥𝔢 𝔄𝔫𝔫𝔞𝔩𝔢𝔫 No. 92, 1924, pp.305–316

Translation by Stephen Bauer-Mengelberg as:

On the building blocks of mathematical logic In [van Heijenoort 67/77], pp.355–366

[Scott 70] Dana S. Scott Lattice Theory, Data Types, and Semantics In Formal Semantics of Programming Languages, Randall Rustin (Editor), pp.65–106

Proc. 2nd Courant Computer Science Symposia, 1970 Prentice Hall, 1972 [Scott 76]

Dana S. Scott Data Types as Lattices To the Memory of Christopher Strachey, 1916–1975 In SIAM Journal on Computing, Vol. 5, No. 3, Sep 1976, pp.522–587

[Sestoft 85] Peter Sestoft The Structure of a Self-Applicable Partial Evaluator In [Ganzinger & Jones 85], pp.236–256

[Sestoft 88] Peter Sestoft Automatic Call Unfolding in a Partial Evaluator In [Bjørner, Ershov & Jones 88], pp.485–506

[Sestoft 88/89] Peter Sestoft Replacing Function Parameters by Global Variables Masters Thesis (“Speciale”), DIKU, University of Copenhagen, Denmark, October 1988

Internet URL: http://www.itu.dk/people/sestoft/mscthesis/

Condensed version:

Joseph E. Stoy (Chairman)

Functional Programming Languages and Computer Architecture (FPCA ’89)

(Proc. of the 4th ACM/IFIP Conference, Imperial College, London, Sep 11–13, 1989) ACM Press, New York, 1989, pp.39–53

[Sestoft 91] Peter Sestoft Analysis and Efficient Implementation of Functional Programs Ph.D. Dissertation, DIKU, University of Copenhagen, Denmark, October 1991

Internet URL:

   ftp://ftp.diku.dk/diku/semantics/papers/D-121.ps.Z

[Sestoft 98] Peter Sestoft Bibliography on partial evaluation and mixed computation On-line BibTEX-format bibliography, last updated: Feb 1998

Internet URL:

   ftp://ftp.diku.dk/diku/semantics/partial-evaluation/

[Shakespeare 1593] William Shakespeare The Taming of the Shrew London: Methuen and Co., 1904

Internet URL: http://books.google.com/books?id=vh4Iau2FIMkC

(Act IV, Scene 1, Line 157). Petruchio’s Country House. Petruchio: A whoreson beetle-headed, flap-ear’d knave!

Internet URL: http://www.gutenberg.org/dirs/etext00/0ws1010.zip

[Shakespeare 1594] William Shakespeare The Tragedy of Romeo and Juliet London: Methuen and Co., 1900

Internet URL: http://books.google.com/books?id=cFzUxZt65lYC

(Act II, Scene 2, Lines 43–44). Capulet’s Orchard. (Balcony) Juliet: What’s in a name? that which we call a rose

By any other name would smell as sweet;

Internet URL: http://www.gutenberg.org/dirs/etext00/0ws1610.zip

[Shakespeare 1603] William Shakespeare The Tragicall Historie of Hamlet, Prince of Denmarke Yale University Press, 1917

Internet URL: http://books.google.com/books?id=GFno8tQ47BoC

(Act II, Scene 2, Line 206). A room in the castle. Polonius. [Aside]: Though this be madness, yet there is method in’t.

(Act III, Scene 1, Lines 56–60). A room in the castle. Hamlet. [Soliloquy]: To be or not to be: that is the question:

Whether ’tis nobler in the mind to suffer The slings and arrows of outrageous fortune, Or to take arms against a sea of troubles, And by opposing end them? To die: to sleep;

Internet URL: http://www.gutenberg.org/dirs/etext00/2ws2610.zip

[Shea & Wilson 76,84] Robert Shea and Robert Anton Wilson The Golden Apple Dell Publishing Co., 1976 In The Illuminatus! Trilogy, pp.438f

Dell Publishing Co., 1984

Excerpt URL: http://www.rawilson.com/illuminatus.shtml#fnord

[Shivers 88] Olin Shivers Control Flow Analysis in Scheme In Proc. ACM PLDI ’88, Jun 1988, pp.164–174

Internet URL:

   http://repository.readscheme.org/ftp/papers/shivers-pldi88.ps.gz
   http://repository.readscheme.org/ftp/papers/shivers-pldi88.pdf

[Shivers 91] Olin Shivers Control-Flow Analysis of Higher-Order Languages or Taming Lambda Ph.D. Dissertation, CMU-CS-91-145, Carnegie-Mellon University, May 1991

Internet URL: http://citeseer.ist.psu.edu/shivers91controlflow.html

[Shoch & Hupp 82] John F. Shoch and Jon A. Hupp The “Worm” Programs—Early Experience with a Distributed Computation In CACM, Vol. 25, No. 3, Mar 1982, pp.172–180

[Siegel 56] Sidney Siegel Nonparametric Statistics: For the Behavioral Sciences (McGraw-Hill Series in Psychology) McGraw-Hill Book Company, Inc., 1956

[Sperber 96] Michael Sperber Self-Applicable Online Partial Evaluation In [Danvy, Glück & Thiemmann 96], pp.465–480

   Internet URL: http://repository.readscheme.org/ftp/papers/sperber-pe96.ps.gz
                 http://repository.readscheme.org/ftp/papers/sperber-pe96.pdf

[Srivastava & Eustace 94] Amitabh Srivastava and Alan Eustace ATOM: A System for Building Customized Program Analysis Tools DEC/WRL 94/2, Digital Equipment Corp., Western Research Lab., Mar 94, also In Proc. ACM PLDI ’94, Jun 1994, pp.196–205 [Steele 77]

Guy Lewis Steele Jr. Debunking the “Expensive Procedure Call” Myth

or, Procedure Call Implementations Considered Harmful or, LAMBDA: The Ultimate GOTO

MIT AI-TR-443, MIT Artificial Intelligence Laboratory, Oct 1977

[Steele 78] Guy Lewis Steele Jr. Rabbit: A Compiler for Scheme (A Study in Compiler Optimization) S.M. Thesis, MIT AI-TR-474, MIT Artificial Intelligence Laboratory, May 1978

[Steele 85/90] Guy L. Steele Jr. Common Lisp: The Language Digital Press, 1984, 1990 (Second Edition)

[Stein 87] Lynn Andrea Stein Delegation is inheritance In Proc. ACM OOPSLA ’87, Oct 1987, pp.138–146, or Also as: ACM SIGPLAN Notices, Vol. 22, No. 12, Dec 1987, pp.138–146

[Stepanov & Lee 95] Alexander Stepano and Meng Lee The Standard Template Library Technical Report 95-11(R.1), HP Laboratories, November 14, 1995

Internet URL: http://www.stepanovpapers.com/STL/DOC.PS
              http://www.stepanovpapers.com/STL/DOC.PDF

[Stoy 77] Joseph E. Stoy Denotational Semantics: The Scott-Strachey Approach to Programming Language Theory (The MIT Press Series in Computer Science) The MIT Press, 1977

[Strachey 73] Christopher Strachey A Mathematical Semantics Which Can Deal with Full Jumps In ^{⟨⟨}Théorie des Algorithmes, des Langages et de la Programmation^{⟩⟩}

(Theories of Algorithms, Languages and Programming)

INRIA, 1973, pp.175–191 [Strachey & Wadsworth 74/2000]

Christopher Strachey and Christopher P. Wadsworth Continuations: A Mathematical Semantics for Handling Full Jumps Technical Monograph PRG-11, Oxford Univiversity Computing Laboratory. Jan 1974

Reprinted (with a forward by Christopher P. Wadsworth) in:

Higher-Order and Symbolic Computation, Vol. 13, Nos. 1–2, 2000, pp.135–152

[Stroustrup 2000] Bjarne Stroustrup The C++ Programming Language [Third Edition] Addison-Wesley, 2000

[Szolovits 92] Peter Szolovits Compilation for fast calculation over pedigrees In Proc. Genetic Analysis Workshop 7, In Cytogenetics and Cell Genetics 59, 1992, pp.136–138 S. Karger Medical and Scientific Publishers

[Takano 91] Akihiko Takano Generalized Partial Computation for a Lazy Functional Language In Proc. ACM SIGPLAN PEPM ’91, New Haven, Connecticut, June 17–19, 1991, pp.1–11 Also as: ACM SIGPLAN Notices, Vol. 26, No. 9, Sep 1991, pp.1–11

[Tanter 2004] Éric Tanter From Metaobject Protocols to Versatile Kernels for Aspect-Oriented Programming Ph.D. Dissertation, Université de Nantes, France & Universidad de Chile, Chile, Nov 2004

Internet URL: http://www.dcc.uchile.cl/~etanter/

[Tarski 55] Alfred Tarski “A lattice-theoretical fixpoint theorem and its applications” In Pacific Journal of Mathematics, Vol. 5, No. 2, June 1955, pp.285–309

Internet URL: http://projecteuclid.org/euclid.pjm/1103044538

[Tomb & Flanagan 2005] Aaron Tomb and Cormac Flanagan Automatic Type Inference via Partial Evaluation In Proc. ACM Int. Conf. on Principles & Practice of Declarative Programming (PPDP ’05)

Lisbon, Portugal, July 11–13, 2005

Internet URL: http://www.cs.ucsc.edu/~atomb/tomb05inference.pdf

[Toole 80,94] John Kennedy Toole A Confederacy of Dunces Louisiana State University Press, 1980 Grove/Atlantic, Inc., 1994

[Traven 27,35] B. Traven 𝔇er Schatz der Sierra Madre (1927)

Translated as:

Treasure of the Sierra Madre (1935) ISBN 0-8090-0160-8

Quotation URL: http://www.darryl.com/badges/badgesbook.gif
Reference URL: http://en.wikipedia.org/wiki/Stinking_badges

[Turing 36–37,65] Alan Mathison Turing “On computable numbers, with an application to the 𝔈𝔫𝔱𝔰𝔠𝔥𝔢𝔦𝔡𝔲𝔫𝔤𝔰𝔭𝔯𝔬𝔟𝔩𝔢𝔪” In Proc. London Mathematical Society, Ser. 2, Vol. 42, 1936–37, pp.230–265

A correction, \underline{Ibid.}, Vol. 43, 1937, pp.544–546

Reproduced in [Davis 65].

[Wahbe, Lucco & Graham 93] Robert Wahbe, Steven Lucco and Susan L. Graham Practical Data Breakpoints: Design and Implementation In Proc. ACM PLDI ’93, Jun 1993, pp.1–12

[Wall 91] David W. Wall Predicting Program Behavior Using Real or Estimated Profiles In Proc. ACM PLDI ’91, Jun 1991, pp.59–70

[Ward & Halstead 90] Stephen A. Ward and Robert H. Halstead, Jr. Computation Structures (The MIT Electrical Engineering and Computer Science Series) The MIT Press, 1990

[Weise, Conybeare, Ruf & Seligman 91] Daniel Weise, Roland Conybeare, Erik Ruf and Scott Seligman Automatic Online Partial Evaluation FUSE-MEMO 91-6, Stanford University Computer Systems Laboratory, Jun 1991 Internet URL:

   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-91-6.ps.gz
   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-91-6.pdf

Also in: [Hughes 91], pp.165–191

See also [Ruf 93, Chapter 3].

[Weise & Ruf 90′] Daniel Weise and Erik Ruf Computing Types During Partial Evaluation Fuse Memo 90-3-revised, CSL-TR-90-441, Stanford Univ. Comp. Systems Lab., Dec 1990

Internet URL:

   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-90-3-REVISED.ps.gz
   http://repository.readscheme.org/ftp/papers/fuse-memos/FUSE-MEMO-90-3-REVISED.pdf

Superseded by [Ruf & Weise 92p].

[Wellstead & Zarrop 91] P. E. Wellstead and M. B. Zarrop Self-tuning Systems: Control and Signal Processing John Wiley & Sons Ltd., 1991

[Whitehead & Russell 10/12/13,25/27,62] Alfred North Whitehead and Bertrand Russell Principia Mathematica (3 Volumes) Cambridge University Press, (1st edition) 1910, 1912, and 1913 Cambridge University Press, (2nd edition) 1925 (Vol. 1), 1927 (Vol. 2 & 3)

Available on-line at the University of Michigan Historical Math Collection:

Vol. 1 URL: http://www.hti.umich.edu/cgi/b/bib/bibperm?q1=AAT3201.0001.001
Vol. 2 URL: http://www.hti.umich.edu/cgi/b/bib/bibperm?q1=AAT3201.0002.001
Vol. 3 URL: http://www.hti.umich.edu/cgi/b/bib/bibperm?q1=AAT3201.0003.001

Abridged as:

Principia Mathematica to *56 Cambridge University Press, 1962 [Widen 86]

Gregory Widen Highlander Directed by Russell Mulcahy Produced by Peter S. Davis, E. C. Monell, William N. Panzer Screenplay by Gregory Widen, Peter Bellwood, Larry Ferguson Story by Gregory Widen Distributed by 20th Century Fox Release date: March 7, 1986 Running time: 116 min.

[van Wijngården 66] Adriaan “Aad” van Wijngården Recursive Definition of Syntax and Semantics In Thomas B. Steel (Editor):

Formal Language Description Languages for Computer Programming, pp.12–24 North-Holland Publising Co., Amsterdam, 1966

[Yamane 67] Taro Yamane Elementary Sampling Theory Prentice-Hall, Inc., 1967

[Young & Hudak 86] Jonathan Young and Paul Hudak Finding Fixpoints on Function Spaces Research Rep. YALEU/DCS/RR-505, Yale University, Dept. of Computer Sci., Dec 1986